OpenStudy (anonymous):

Subtract and simplify: (3x2 – 6x + 7) – (x2 – 6x + 4) A. 2x2 + 6x + 11 B. 2x2 + 3 C. 4x2 – 6x + 3 D. 2x2 + 6x + 3

OpenStudy (anonymous):

@hoblos

OpenStudy (anonymous):

i picked A

OpenStudy (hoblos):

can show me how you simplified it ?

OpenStudy (anonymous):

im assuming its wrong

OpenStudy (hoblos):

yes.. actually it is (3x2 – 6x + 7) – (x2 – 6x + 4) = 3x2 – 6x + 7 – x2 + 6x - 4 = (3-1)x2 + (-6+6)x +7 - 4 = 2x2 + 3

OpenStudy (anonymous):

i will review it thank you

OpenStudy (hoblos):

glad to help :)